OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 * x + 5 * 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

first step is to eliminate the ( ) can you distribute the 5 in this part? 5(8+7)

OpenStudy (anonymous):

Sure! 5*8+7 is that right or is that Order of Operations. I can't remember.

OpenStudy (anonymous):

use pemdas.. order of operations

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

5(8+7) = (5)8 + (5)7 so now we have (5)8 + (5)7 = 5 * x + 5 * 7 what do you think we should do next?

OpenStudy (anonymous):

I am trying to think, but I can't remember.

OpenStudy (anonymous):

lets 'isolate' the x term. so on the right hand side, we have +5*7, how can we get that to the other side of the equal sign?

OpenStudy (anonymous):

hmmmm...

OpenStudy (anonymous):

subtract -(5*7) from both sides. what does your equation look like now?

OpenStudy (anonymous):

You may think I am stupid, but I seriously don't know. :(

OpenStudy (anonymous):

it's ok!

OpenStudy (anonymous):

youre not dumb, you just haven't learned yet.

OpenStudy (anonymous):